# Fast Robust Kernel Regression through Sign Gradient Descent with Early Stopping

### v6

@@ Appendix C Proofs
Proof of Lemma 4.
@@
$\bm{0}$ $=\frac{\partial}{\partial\bm{\hat{f}^{{}^{+}}}}\left(\frac{1}{2}\left\|\bm{y^{{}^{+}}}-\bm{\hat{f}^{{}^{+}}}\right\|_{\bm{K^{**}}}^{2}+\frac{\lambda}{2}\|\bm{\hat{f}^{{}^{+}}}\|^{2}_{2}\right)=-\bm{K^{**}}\bm{y^{{}^{+}}}+\bm{K^{**}}\bm{\hat{f}^{{}^{+}}}+\lambda\bm{\hat{f}^{{}^{+}}}$
@@
Proof of Lemma 5 with momentum and Nesterov accelerated gradient.
Analogous to the case of $\bm{\hat{\alpha}}$ in the proof of the remark of Lemma 2, for momentum and Nesterov accelerated gradient, Equation 19 generalizes into
@@
Solving the differential equations in the same way as in Lemma 5, we obtain
@@
$\text{sign}\left(-\frac{\partial}{\partial\bm{f^{{}^{+}}}(t)}\left(\left\|\bm{y^{{}^{+}}}-\bm{f^{{}^{+}}}(t)\right\|_{\bm{K^{**}}}^{2}\right)\right)=\text{sign}\left(\bm{K^{**}}(\bm{y^{{}^{+}}}-\bm{f^{{}^{+}}}(t))\right)$
$=\text{sign}\left(\begin{bmatrix}\vrule\\ k_{ii}(y^{{}^{+}}_{i}-f^{{}^{+}}_{i}(t))\\ \vrule\end{bmatrix}\right)=\text{sign}\left(\begin{bmatrix}[k_{ii}(y_{i}-f_{i}(t))]_{i=1}^{n}\\ \bm{0}\end{bmatrix}\right)$
$=\begin{bmatrix}[\text{sign}(k_{ii}(y_{i}-f_{i}(t)))]_{i=1}^{n}\\ \bm{0}\end{bmatrix},$
where $[\text{sign}(k_{ii}(y_{i}-f_{i}(t)))]_{i=1}^{n}\in\mathbb{R}^{n}$ and $\bm{0}\in\mathbb{R}^{n^{*}}$. Since element $i$ in the vector depends on $\bm{f^{{}^{+}}}$ only through $f^{{}^{+}}_{i}$, and since $k_{ii}>0$,
$\frac{\partial\bm{f^{{}^{+}}}(t)}{\partial t}=\text{sign}\left(-\frac{\partial}{\partial\bm{f^{{}^{+}}}(t)}\left(\left\|\bm{y^{{}^{+}}}-\bm{f^{{}^{+}}}(t)\right\|_{\bm{K^{**}}}^{2}\right)\right)$
$\iff$ $\frac{\partial f^{{}^{+}}_{i}(t)}{\partial t}=\begin{cases}\text{sign}(k_{ii}(y_{i}-f_{i}(t)))&\text{if }i\leq n\\ 0&\text{if }i>n\end{cases}=\begin{cases}1&\text{if }i\leq n\text{ and }y_{i}>f^{{}^{+}}_{i}(t)\\ -1&\text{if }i\leq n\text{ and }y_{i}<f^{{}^{+}}_{i}(t)\\ 0&\text{if }i>n\text{ or }y_{i}=f^{{}^{+}}_{i}(t).\end{cases}$
Thus, since $f^{{}^{+}}_{i}(0)=0$, $f^{{}^{+}}_{i}(t)=0$, for $i>n$. For $i\leq n$, $f^{{}^{+}}_{i}(t)=\pm t$, depending on the sign of $y_{i}$. Once $f^{{}^{+}}_{i}=y_{i}$ (and $\text{sign}(k_{ii}(y_{i}-f^{{}^{+}}_{i})=0$), then $f^{{}^{+}}_{i}=y$:

### v7

@@ Appendix D Proofs
Proof of Proposition 5.3.
For KGD, let $L_{2}(\bm{\hat{\alpha}}):=\|\bm{y}-\bm{K}\bm{\hat{\alpha}}\|_{2}^{2}$. Then $L_{2}^{\prime}(\bm{\hat{\alpha}})=-\bm{K}(\bm{y}-\bm{K}\bm{\alpha})$, $L_{2}^{\prime\prime}(\bm{\hat{\alpha}})=\bm{K}^{2}$ and higher order derivatives are zero. With $\bm{\Delta\hat{\alpha}}=\eta\cdot(\bm{y}-\bm{K}\bm{\hat{\alpha}})$, according to Taylor’s theorem
$L_{2}(\bm{\hat{\alpha}}_{k+1})-L_{2}(\bm{\hat{\alpha}}_{k})=L_{2}(\bm{\hat{\alpha}}_{k}+\bm{\Delta\hat{\alpha}}_{k})-L_{2}(\bm{\hat{\alpha}}_{k})$
$=L_{2}(\bm{\hat{\alpha}}_{k})+L_{2}^{\prime}(\bm{\hat{\alpha}}_{k})^{\top}\bm{\Delta\hat{\alpha}}+\frac{1}{2}\bm{\Delta\hat{\alpha}}^{\top}L_{2}^{\prime\prime}(\bm{\hat{\alpha}}_{k})\bm{\Delta\hat{\alpha}}-L_{2}(\bm{\hat{\alpha}}_{k})$
$=-\eta\cdot(\bm{y}-\bm{K}\bm{\alpha})^{\top}\bm{K}(\bm{y}-\bm{K}\bm{\hat{\alpha}})+\frac{\eta^{2}}{2}(\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k})^{\top}\bm{K}^{2}(\bm{y}-\bm{K}\bm{\hat{\alpha}})$
$\leq-\eta\cdot(\bm{y}-\bm{K}\bm{\alpha})^{\top}\bm{K}(\bm{y}-\bm{K}\bm{\hat{\alpha}})+\frac{\eta^{2}}{2}s_{\max}(\bm{K})(\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k})^{\top}\bm{K}(\bm{y}-\bm{K}\bm{\hat{\alpha}}).$
Factoring out $\eta\cdot(\bm{y}-\bm{K}\bm{\alpha})^{\top}\bm{K}(\bm{y}-\bm{K}\bm{\hat{\alpha}})$, and using that $(\bm{y}-\bm{K}\bm{\alpha})^{\top}\bm{K}(\bm{y}-\bm{K}\bm{\hat{\alpha}})\geq 0$, since $\bm{K}$ is positive semi-definite, we see that $L_{2}(\bm{\hat{\alpha}}_{k+1})-L_{2}(\bm{\hat{\alpha}}_{k})\leq 0$ if $\eta\leq\frac{2}{s_{\max}(\bm{K})}$.
For KSGD, $L_{1}(\bm{\hat{\alpha}}):=\|\bm{y}-\bm{K}\bm{\hat{\alpha}}\|_{1}=\text{sign}(\bm{y}-\bm{K}\bm{\hat{\alpha}})^{\top}(\bm{y}-\bm{K}\bm{\hat{\alpha}})$. Then $L_{1}^{\prime}(\bm{\hat{\alpha}})=-\bm{K}\cdot\text{sign}(\bm{y}-\bm{K}\bm{\alpha})$ and higher order derivatives are zero. With $\bm{\Delta\hat{\alpha}}=\eta\cdot\text{sign}(\bm{y}-\bm{K}\bm{\hat{\alpha}})$, according to Taylor’s theorem, for $\eta\leq\min_{1\leq i\leq n}\left(\left|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\right|_{i}\right)$
$L_{1}(\bm{\hat{\alpha}}_{k+1})-L_{1}(\bm{\hat{\alpha}}_{k})$ $=L_{1}(\bm{\hat{\alpha}}_{k}+\bm{\Delta\hat{\alpha}}_{k})-L_{1}(\bm{\hat{\alpha}}_{k})=L_{1}(\bm{\hat{\alpha}}_{k})+L^{\prime}(\bm{\hat{\alpha}}_{k})^{\top}\bm{\Delta\hat{\alpha}}$
$=-\eta\cdot\text{sign}(\bm{y}-\bm{K}\bm{\alpha})^{\top}\bm{K}\cdot\text{sign}(\bm{y}-\bm{K}\bm{\hat{\alpha}})\leq 0.$
For KCD, let $\mathbb{I}_{\infty}(\bm{v})$ be a vector denoting the sign of the largest (absolute) value in $\bm{v}$, such that
$\mathbb{I}_{\infty}(\bm{v})_{d}=\begin{cases}\text{sign}(v_{d})&\text{ if }d=\operatorname*{arg\,max}_{d^{\prime}}|v_{d^{\prime}}|\\ 0&\text{ else,}\end{cases}$
so that $\|\bm{v}\|_{\infty}=\mathbb{I}_{\infty}(\bm{v})^{\top}\bm{v}$. Furthermore, let $L_{\infty}(\bm{\hat{\alpha}}):=\|\bm{y}-\bm{K}\bm{\hat{\alpha}}\|_{\infty}=\mathbb{I}_{\infty}(\bm{y}-\bm{K}\bm{\hat{\alpha}})^{\top}(\bm{y}-\bm{K}\bm{\hat{\alpha}})$. Then $L_{\infty}^{\prime}(\bm{\hat{\alpha}})=-\bm{K}\cdot\mathbb{I}_{\infty}(\bm{y}-\bm{K}\bm{\alpha})$ and higher order derivatives are zero. With $\bm{\Delta\hat{\alpha}}=\eta\cdot\mathbb{I}_{\infty}(\bm{y}-\bm{K}\bm{\hat{\alpha}})$, according to Taylor’s theorem, for $\eta\leq\min_{1\leq i\leq n}\left(\left|\bm{y}-\bm{K}\bm{\hat{\alpha}}_{k}\right|_{i}\right)$
$L_{\infty}(\bm{\hat{\alpha}}_{k+1})-L_{\infty}(\bm{\hat{\alpha}}_{k})$ $=L_{\infty}(\bm{\hat{\alpha}}_{k}+\bm{\Delta\hat{\alpha}}_{k})-L_{\infty}(\bm{\hat{\alpha}}_{k})=L_{\infty}(\bm{\hat{\alpha}}_{k})+L^{\prime}(\bm{\hat{\alpha}}_{k})^{\top}\bm{\Delta\hat{\alpha}}$
$=-\eta\cdot\mathbb{I}_{\infty}(\bm{y}-\bm{K}\bm{\alpha})^{\top}\bm{K}\cdot\mathbb{I}_{\infty}(\bm{y}-\bm{K}\bm{\hat{\alpha}})\leq 0.$
Proof of Proposition A.1.
@@
$\bm{0}$ $=\frac{\partial}{\partial\bm{\hat{f}^{{}^{+}}}}\left(\frac{1}{2}\left\|\bm{y^{{}^{+}}}-\bm{\hat{f}^{{}^{+}}}\right\|_{\bm{K^{**}}}^{2}+\frac{\lambda}{2}\|\bm{\hat{f}^{{}^{+}}}\|^{2}_{2}\right)$
$=-\bm{K^{**}}\bm{y^{{}^{+}}}+\bm{K^{**}}\bm{\hat{f}^{{}^{+}}}+\lambda\bm{\hat{f}^{{}^{+}}}$
@@
Proof of Equation 19 with momentum and Nesterov accelerated gradient.
Analogous to the case of $\bm{\hat{\alpha}}$ in the proof of the Remark 4.2, for momentum and Nesterov accelerated gradient, Equation 18 generalizes into
@@
Solving the differential equations in the same way as for Equation 19, we obtain
@@
$\text{sign}\left(-\frac{\partial}{\partial\bm{\hat{f}^{{}^{+}}}(t)}\left(\left\|\bm{y^{{}^{+}}}-\bm{\hat{f}^{{}^{+}}}(t)\right\|_{\bm{K^{**}}}^{2}\right)\right)=\text{sign}\left(\bm{K^{**}}(\bm{y^{{}^{+}}}-\bm{\hat{f}^{{}^{+}}}(t))\right)$
$=\text{sign}\left(\begin{bmatrix}\vrule\\ k_{ii}(y^{{}^{+}}_{i}-\hat{f}^{{}^{+}}_{i}(t))\\ \vrule\end{bmatrix}\right)=\text{sign}\left(\begin{bmatrix}[k_{ii}(y_{i}-\hat{f}_{i}(t))]_{i=1}^{n}\\ \bm{0}\end{bmatrix}\right)$
$=\begin{bmatrix}[\text{sign}(k_{ii}(y_{i}-\hat{f}_{i}(t)))]_{i=1}^{n}\\ \bm{0}\end{bmatrix},$
where $[\text{sign}(k_{ii}(y_{i}-\hat{f}_{i}(t)))]_{i=1}^{n}\in\mathbb{R}^{n}$ and $\bm{0}\in\mathbb{R}^{n^{*}}$. Since element $i$ in the vector depends on $\bm{\hat{f}^{{}^{+}}}$ only through $\hat{f}^{{}^{+}}_{i}$, and since $k_{ii}>0$,
$\frac{\partial\bm{\hat{f}^{{}^{+}}}(t)}{\partial t}$ $=\text{sign}\left(-\frac{\partial}{\partial\bm{\hat{f}^{{}^{+}}}(t)}\left(\left\|\bm{y^{{}^{+}}}-\bm{\hat{f}^{{}^{+}}}(t)\right\|_{\bm{K^{**}}}^{2}\right)\right)$
$\iff\frac{\partial\hat{f}^{{}^{+}}_{i}(t)}{\partial t}$ =$=\begin{cases}\text{sign}(k_{ii}(y_{i}-\hat{f}_{i}(t)))&\text{if }i\leq n\\ 0&\text{if }i>n\end{cases}$
=$=\begin{cases}1&\text{if }i\leq n\text{ and }y_{i}>\hat{f}^{{}^{+}}_{i}(t)\\ -1&\text{if }i\leq n\text{ and }y_{i}<\hat{f}^{{}^{+}}_{i}(t)\\ 0&\text{if }i>n\text{ or }y_{i}=\hat{f}^{{}^{+}}_{i}(t).\end{cases}$
Thus, since $\hat{f}^{{}^{+}}_{i}(0)=0$, $\hat{f}^{{}^{+}}_{i}(t)=0$, for $i>n$. For $i\leq n$, $\hat{f}^{{}^{+}}_{i}(t)=\pm t$, depending on the sign of $y_{i}$. Once $\hat{f}^{{}^{+}}_{i}=y_{i}$ (and $\text{sign}(k_{ii}(y_{i}-\hat{f}^{{}^{+}}_{i})=0$), then $\hat{f}^{{}^{+}}_{i}=y$: